# Accelerated Ostwald ripening by chemical activity

### v1

@@ Abstract
Phase separation of biomolecular condensates is ubiquitous in living cells, promoting colocalization of enzymes and their substrates as well as achieving membrane-free compartmentalization. Energy-consuming processes are routinely used to regulate biocondensate growth by opposing the thermodynamic tendency toward coarsening. At the same time, cells often use energy to instead accelerate thermodynamic processes. Here, we theoretically explore the possibility of utilizing chemical reactions to accelerate biocondensate coarsening. We combine Lifshitz-Slyozov theory with a reaction-diffusion approach, wherein particles interconvert between phase-separating and inert forms. We find that mass conservation restricts the volume growth to be linear in time (as in the passive case) despite activity, though if reactions are restricted to occur only outside droplets, the rate of Ostwald ripening can be increased by an arbitrarily large factor. Our acceleration theory is quantitatively supported by recent experiments on ripening in the presence of fueled interconversion reactions, under precisely the predicted conditions. We posit that the ability to induce rapid biocondensate coarsening can be advantageous in synthetic-biological contexts as a regulator of metabolic channeling.
I Introduction
@@
Ostwald ripening refers to the thermodynamic process of coarsening of solute-rich phase-separated droplets. As a result of surface tension, smaller droplets shrink due to their larger Laplace pressure whereas larger droplet grow via the diffusion of solute. Average droplet size may also increase due to coalescence events. These processes continue in principle until phase-separation completes. However, a vastly richer physics emerges when activity and energy consumption are introduced. One well-studied outcome is “size control” [1, 2, 3, 4, 5], whereby active processes prevent unbounded coarsening. Size control can be achieved via nonequilibrium reactions that interconvert the solute between phase-separating and inert forms. Additional possible outcomes of activity include arrested ripening [6] (where small droplets do not shrink), creation of Turing patterns [7, 8] (where the dense phase is destabilized by the coupling between thermodynamics and the reactions), and reversal of Ostwald ripening [9] (where dilute bubbles form inside the dense phase). Further complexities such as nonlocal elasticity [10, 11, 12, 13] or charge asymmetries [14] extend the list of possible outcomes. In contrast to these effects which oppose Ostwald ripening, it was recently demonstrated experimentally that fueled interconversion reactions in organochemical emulsions can subtantially accelerate ripening [15]. How do these interconversion reactions lead to such a different outcome?
@@
One motivation to investigate accelerated Ostwald ripening is its potential biological relevance. Phase separation of biomolecules to form biomolecular condensates (also termed “membraneless organelles”) is pervasive in cells, enabling the compartmentalization of biochemical processes [16, 17]. Biocondensate size can be crucial for function [18, 19, 20, 21, 22]. For example, consider the algal pyrenoid [23] — a condensate composed of Rubisco (a carbon-fixating enzyme) and EPYC1 (a “sticky” linker protein that drives phase separation) — which is dynamically regulated through energy-consuming phosphorylation cycles. During cell division, phosphorylation renders EPYC1 less sticky, breaking the pyrenoid into small droplets via a size-control mechanism, thus favoring equipartition of Rubisco between the daughter cells. After division, dephosphorylation allows recoarsening into a single pyrenoid in each daughter cell. For cases like the pyrenoid where a single condensate, or a small number of large condensates, is advantageous, might cells use energy to accelerate Ostwald ripening, and how might this be achieved?
@@
Figure 1: Illustration of the model. (a) Passive system: A slightly supersaturated, monodisperse suspension of “sticky” particles. Spontaneously-formed dense-phase nuclei will coarsen at the expense of the material available in the surrounding dilute phase. Once the supersaturation becomes small, Ostwald ripening commences — while larger droplets keep growing, smaller droplets, whose Laplace pressure is higher, begin shrinking. (b) Active system: Motivated by the active ripening of the algal pyrenoid , we introduce kinases and phosphatases to the above suspension. The former phosphorylate the sticky proteins, rendering them “nonsticky” (inert), and the latter dephosphorylate nonsticky proteins returning them to the sticky form. In our study, we investigate whether these active processes can allow for faster Ostwald ripening.

### v2

@@ Abstract
Phase separation of biomolecular condensates promotes membrane-free compartmentalization in cells. The dynamics of these biocondensates is routinely regulated by energy-consuming processes. Here, we devise a theory pinpointing how active chemical reactions, interconverting molecules between phase-separating and inert forms, can drive faster condensate coarsening. We find that mass conservation limits droplet volume growth to being linear in time regardless of activity, resembling the passive Lifshitz-Slyozov law. However, if reactions are restricted to occur only outside droplets, the rate of Ostwald ripening can be increased by an arbitrarily large factor. Our theory is quantitatively supported by recent experiments on ripening in the presence of fueled interconversion reactions, under precisely the predicted conditions. We posit that the ability to induce rapid biocondensate coarsening can be advantageous in synthetic-biological contexts, e.g., as a regulator of metabolic channeling.
@@
Ostwald ripening refers to the thermodynamic coarsening process of solute-rich phase-separated droplets as a result of surface tension. Smaller droplets shrink due to their larger Laplace pressure whereas larger droplet grow via the diffusion of solute. Average droplet size may also increase due to coalescence events. These processes continue in principle until phase-separation completes. However, a vastly richer physics emerges when activity and energy consumption are introduced. One well-studied outcome is “size control” [1, 2, 3, 4, 5], whereby active processes prevent unbounded coarsening. Size control can be achieved via nonequilibrium reactions that interconvert the solute between phase-separating and inert forms. Additional possible outcomes of activity include arrested ripening [6] (where small droplets do not shrink), emergence of Turing patterns [7, 8] (where the dense phase is destabilized by the coupling between thermodynamics and reactions), and reversal of Ostwald ripening [9] (where dilute bubbles form inside the dense phase). Further complexities such as elasticity [10, 11, 12, 13] or charge asymmetry [14] extend the list of possible outcomes. In contrast to these effects which oppose Ostwald ripening, it was recently demonstrated experimentally that fueled interconversion reactions in organochemical emulsions can subtantially accelerate ripening [15]. How do these same interconversion reactions lead to such a different outcome?
@@
One motivation to investigate accelerated Ostwald ripening is its potential biological relevance. Phase separation of biomolecules to form biomolecular condensates (also termed “membraneless organelles”) is pervasive in cells, enabling the compartmentalization of biochemical processes [16, 17]. Biocondensate size can be crucial for function [18, 19, 20, 21, 22]. For example, consider the algal pyrenoid [23] — a condensate composed of Rubisco (a carbon-fixating enzyme) and EPYC1 (a “sticky” linker protein that drives phase separation) — which is dynamically regulated through energy-consuming phosphorylation cycles. During cell division, phosphorylation renders EPYC1 less sticky, breaking the pyrenoid into small droplets via a size-control mechanism, thus favoring equipartition of Rubisco between the daughter cells. After division, dephosphorylation allows recoarsening into a single pyrenoid in each daughter cell. For cases like the pyrenoid where a single condensate is advantageous, could cells use energy to accelerate Ostwald ripening, and how might this be achieved?
@@
Figure 1: Illustration of the model. (a) Passive system: A supersaturated monodisperse suspension of “sticky” particles. Spontaneously-formed dense-phase nuclei will coarsen at the expense of the material available in the surrounding dilute phase. Once the supersaturation becomes small, Ostwald ripening commences — larger droplets keep growing while smaller droplets, whose Laplace pressure is higher, begin shrinking. (b) Active system: Motivated by the active ripening of the algal pyrenoid [23], we introduce kinases and phosphatases to the above suspension. The former phosphorylate the sticky proteins, rendering them “nonsticky” (inert), and the latter dephosphorylate nonsticky proteins returning them to the sticky form. In our study, we investigate whether these active processes can allow for faster Ostwald ripening.